OpenStudy (anonymous):

solve the triangle using either the Law of Sines or the Law of Cosines.Round your answers to the nearest tenth. a = 12, b = 14, c = 21

OpenStudy (anonymous):

solve for the largest angle first, which will be the angle opposite the longest side \(c\)

OpenStudy (anonymous):

\[\cos(C)=\frac{a^2+b^2-c^2}{2ab}\] in your case \[\cos(C)=\frac{12^2+14^2-21^2}{2\times 12\times 14}=-\frac{101}{336}\]

OpenStudy (anonymous):

arccosine of \(-\frac{101}{336}\) is \(107.5^o\) and that is your largest angle now go to the law of cosines

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I would use the sine law to solve the rest of the triangles. Far easier.

OpenStudy (anonymous):

ooops man i am screwing up all night that is what i meant "then go to the law of sines" much easier

OpenStudy (anonymous):

@satellite73 : It's fine: I said the square root of 2 was 1 once.

OpenStudy (anonymous):

yeah an sometimes \((\frac{1}{2})^2=1\) as well

OpenStudy (anonymous):

Hehe...